OpenStudy (anonymous):

f f(x) = x + 4, find f(3). 1 7 −1 12

OpenStudy (anonymous):

I keep getting 6.

hero (hero):

Hint: f(x) = x + 4 f(3) = 3 + 4

OpenStudy (anonymous):

I'm still getting 6?

OpenStudy (anonymous):

(3)+4=?

hero (hero):

How? What is 3 + 4 ?

OpenStudy (anonymous):

you just have to put 3 in the function

OpenStudy (anonymous):

Oh, 7. >.< I'm so slow. I'm sorry :l

OpenStudy (anonymous):

Alright so I did t right just added wrong, thanks!